40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2642     Entry time: Fri Feb 26 01:00:07 2010
Author: Jenne 
Type: Update 
Category: COC 
Subject: Suspension Progress 

This is going to be a laundry list of the mile markers achieved so far:

* Guiderod and wire standoff glued to each ITMX and ITMY

* Magnets glued to dumbbells (4 sets done now).  ITMX has 244 +- 3 Gauss, ITMY has 255 +- 3 Gauss.  The 2 sets for SRM and PRM are 255 +- 3 G and 264 +- 3 G.  I don't know which set will go with which optic yet.

* Magnets glued to ITMX.  There were some complications removing the optic from the magnet gluing fixture.  The way the optic is left with the glue to dry overnight is with "pickle picker" type grippers holding the magnets to the optic.  After the epoxy had cured, Kiwamu and I took the grippers off, in preparation to remove the optic from the fixture.  The side magnet (thankfully the side where we won't have an OSEM) and dumbbell assembly snapped off.  Also, on the UL magnet, the magnet came off of the dumbbell (the dumbbell was still glued to the glass).  We left the optic in the fixture (to maintain the original alignment), and used one of the grippers to glue the magnet back to the UL dumbbell.  The gripper in the fixture has very little slop in where it places the magnet/dumbbell, so the magnet was reglued with very good axial alignment.  Since after the side magnet+dumbbell came off the glass, the 2 broke apart, we did not glue them back on to the optic.  They were reattached, so that we can in the future put the extra side magnet on, but I don't think that will be necessary, since we already know which side the OSEM will be on.

* Magnets glued to ITMY.  This happened today, so it's drying overnight.  Hopefully the grippers won't be sticky and jerky like last time when we were removing them from the fixture, so hopefully we won't lose any magnets when I take the optic out of the fixture.

* ITMX has been placed in its suspension cage.  The first step, before getting out the wire, is to set the optic on the bottom EQ stops, and get the correct height and get the optic leveled, to make things easier once the wire is in place.  Koji and I did this step, and then we clamped all of the EQ stops in place to leave it for the night.

* The HeNe laser has been leveled, to a beam height of 5.5inches, in preparation for the final leveling of the optics, beginning tomorrow.  The QPD with the XY decoder is also in place at the 5.5 inch height for the op lev readout.  The game plan is to leave this set up for the entire time that we're hanging optics.  This is kind of a pain to set up, but now that it's there, it can stay out of the way huddled on the side of the flow bench table, ready for whenever we get the ETMs in, and the recoated PRM. 

* Koji and Steve got the ITMX OSEMs from in the vacuum, and they're ready for the hanging and balancing of the optic tomorrow.  Also, they got out the satellite box, and ran the crazy-long cable to control the OSEMs while they're on the flow bench in the clean room.

 

Koji and I discovered a problem with the small EQ stops, which will be used in all of the SOS suspensions for the bottom EQ stops.  They're too big.  :(  The original document (D970312-A-D) describing the size for these screws was drawn in 1997, and it calls for 4-40 screws.  The updated drawing, from 2000 (D970312-B-D) calls for 6-32 screws.  I naively trusted that updated meant updated, and ordered and prepared 6-32 screws for the bottom EQ stops for all of the SOSes.  Unfortunately, the suspension towers that we have are tapped for 4-40.  Thumbs down to that.  We have a bunch of vented 4-40 screws in the clean room cabinets, which I can drill, and have Bob rebake, so that Zach and Mott can make viton inserts for them, but that will be a future enhancement.  For tonight, Koji and I put in bare vented 4-40 screws from the clean room supply of pre-baked screws.  This is consistent with the optics in our chambers having bare screws for the bottom EQ stops, although it might be nicer to have cushy viton for emergencies when the wire might snap.  The real moral of this story is: don't trust the drawings.  They're good for guidelines, but I should have confirmed that everything fit and was the correct size.

ELOG V3.1.3-